OpenStudy (anonymous):

6^3 = 2^n + 3^n solve for n

OpenStudy (anonymous):

\[\left( \begin{array}{cc} n=1 & 2^n+3^n= 5 \\ n= 2 & 2^n+3^n= 13 \\ n= 3 & 2^n+3^n= 35 \\ n= 4 & 2^n+3^n= 97 \\ n= 5 & 2^n+3^n= 275 \\ n= 6 & 2^n+3^n= 793 \\ \end{array} \right) \] and \[ 6^3=216\] So there is no such n